# Dual LoRA: Enhancing LoRA with Magnitude and Direction Updates

### v1

@@ 4.2 Neural Language Understanding (NLU)
Datasets and baseline models. We evaluate our method on a widely used natural language understanding dataset GLUE. It consists of eight different datasets includes MNLI Williams et al. (2017), SST-2 Socher et al. (2013), MRPC Dolan and Brockett (2005), CoLA Warstadt (2019), QNLI Rajpurkar et al. (2018), QQP, RTE, and STS-B Cer et al. (2017). The diversity makes the GELU benchmark a robust dataset for evaluating LLMs on NLU tasks.
For the baseline models, we use RoBERTa base/large Liu (2019) and DeBERTa XXL He et al. (2020) as pretrained baseline models from the HuggingFace Transformers library Wolf et al. (2020). Similarly to LoRA, we initialize the model to the LoRA-adapted MNLI checkpoint for MRPC, RTE, and STSB rather than the pre-trained baseline model. All other training parameters are the same as LoRA except for the learning rate.
@@ 4.3 Ablation Study
Dealing with the sign function. In the previous section and experiments, we use the straight-through estimator (STE) method Bengio et al. (2013) to compute the gradient of the sign function. Note that the sign function is a standard operator in the area of binary neural networks (BNNs), and there are many studies on dealing with the forward and backward passes of the sign function. For example, XNOR-Net Rastegari et al. (2016) scales the weights after binarized:
@@
where $\textbf{E}_{F}(|x|)$ is the mean of the absolute value of each output channel of weights. Dorefa-Net Zhou et al. (2016) uses a constant scalar to scale all of the weights instead of doing channel-wise scaling:
@@
STE (ours) 1.64 78.9
XNOR-Net 1.64 77.9
Dorefa-Net 1.64 78.1
The experimental results of using different methods to deal with the sign function are shown in Tab. 4.3. The original STE method performs best among different methods. This conclusion is different from that in BNNs. We analyze that this is because both XNOR-Net Rastegari et al. (2016) and Dorefa-Net Zhou et al. (2016) modify the forward pass of sign function by adding per-channel scales or a constant scale, which contaminate the direction group and make it unable to focus on giving the correct binary outputs.
@@
Specifically, we keep the total trainable parameters unchanged by setting $r_{1}+r_{2}=2r$, and adjust the ratio of parameters in the magnitude group and direction group, which are controlled by $r_{1}$ and $r_{2}$, respectively. We conduct experiments on the commonsense reasoning dataset with LLaMA3-8B as our baseline model and set 15 different ratios. Other training hyper-parameters are kept the same as in the previous experiments. The results are shown in Fig. 2. We can see that the proposed Dual LoRA can consistently outperform LoRA and DoRA when having a roughly balanced parameter ratio between the magnitude group and direction group (from 25% to 70%), which shows the robustness of our method.
@@ Appendix A Experiments on Neural Language Generation (NLG)
Datasets. We conduct experiments on E2E NLG Challenge Novikova et al. (2017), DART Nan et al. (2020) and WebNLG Gardent et al. (2017) datasets. The E2E dataset consists of approximately 42,000 training data, 4,600 validation data, and 4,600 test data. Each sample is composed of a sequence of slot-value pairs $(x,y)$ and a corresponding natural language reference text. DART is an open-domain data-to-text dataset with around 82,000 samples, each sample is structured as a sequence of entity-relation-entity triple. WebNLG has a total of 22,000 examples from 14 different categories, each sample is structured as a sequence of subject-property-object triple.
Results. As in previous experiments, we reduce our rank to half of the other LoRA-based methods to ensure the same number of trainable parameters. All other hyper-parameters are the same as LoRA, except that we tune the learning rate. As the results shown in Tab. 4 and Tab. 5, Dual LoRA can consistently outperform all other competitors on three different datasets with baseline model GPT2_M and GPT2_L.
@@ B.1 Different forms of the update matrix
Table 5: The results of the proposed Dual LoRA and other competitors with GPT2_M and GPT2_L on DART and WebNLG datasets. The up arrow indicates that the higher is better, and the down arrow indicates that the lower is better.
Model Methods Trainable DART WebNLG
[…]
GPT2_M LoRA ($r=4$) 0.35 47.4 0.36 0.47 55.0 0.37 0.39
LoRA ($r=8$) 0.70 47.5 0.36 0.47 55.6 0.38 0.39
LoRA+ ($r=8$) 0.70 47.6 0.36 0.47 56.1 0.38 0.39
DoRA ($r=8$) 0.71 47.0 0.36 0.48 53.5 0.36 0.40
Dual LoRA ($r=4$) 0.70 48.3 0.36 0.47 56.6 0.38 0.38
GPT2_L LoRA ($r=4$) 0.35 47.7 0.36 0.47 57.1 0.37 0.38
LoRA ($r=8$) 0.70 47.5 0.36 0.47 57.6 0.38 0.38
LoRA+ ($r=8$) 0.70 47.7 0.36 0.47 57.5 0.38 0.38
DoRA ($r=8$) 0.71 47.2 0.36 0.47 57.4 0.39 0.38
Dual LoRA ($r=4$) 0.70 48.4 0.36 0.47 57.7 0.39 0.38
@@
Dual LoRA ($r=32$) 1.64 70.0 83.2 79.5 87.3 83.0 81.6 65.2 81.0 78.9
Setting 1 ($r=32$) 1.64 70.3 82.9 78.8 85.4 83.3 82.1 65.2 79.4 78.4
Setting 2 ($r=32$) 1.64 70.6 80.8 78.9 84.7 81.8 80.4 65.2 78.8 77.7
Setting 3 ($r=32$) 0.84 60.4 0.2 1.7 0.2 0.2 0.5 0 0 7.9
Setting 3 ($r=64$) 1.64 36.3 36.2 6.6 0 14.3 4.9 11.5 0 13.7
@@
ReLU (ours) 70.0 83.2 79.5 87.3 83.0 81.6 65.2 81.0 78.9
Abs 68.6 82.9 77.9 74.3 74.0 81.6 65.8 79.6 75.6
Sigmoid 66.5 80.3 78.5 77.8 79.4 75.4 60.9 76.6 74.4

### v3

@@ 4.2 Neural Language Understanding (NLU)
Datasets and baseline models. We evaluate our method on a widely used natural language understanding dataset GLUE. It consists of eight different datasets includes MNLI (Williams et al., 2017), SST-2 (Socher et al., 2013), MRPC (Dolan and Brockett, 2005), CoLA (Warstadt, 2019), QNLI (Rajpurkar et al., 2018), QQP, RTE, and STS-B (Cer et al., 2017). The diversity makes the GELU benchmark a robust dataset for evaluating LLMs on NLU tasks.
For the baseline models, we use RoBERTa base/large (Liu, 2019) and DeBERTa XXL (He et al., 2020) as pretrained baseline models from the HuggingFace Transformers library (Wolf et al., 2020). Similarly to LoRA, we initialize the model to the LoRA-adapted MNLI checkpoint for MRPC, RTE, and STSB rather than the pre-trained baseline model. All other training parameters are the same as LoRA except for the learning rate.
@@ 4.3 Ablation Study
Dealing with the sign function. In the previous section and experiments, we use the straight-through estimator (STE) method (Bengio et al., 2013) to compute the gradient of the sign function. Note that the sign function is a standard operator in the area of binary neural networks (BNNs), and there are many studies on dealing with the forward and backward passes of the sign function. For example, XNOR-Net (Rastegari et al., 2016) scales the weights after binarized:
@@
where $\textbf{E}_{F}(|x|)$ is the mean of the absolute value of each output channel of weights. Dorefa-Net (Zhou et al., 2016) uses a constant scalar to scale all of the weights instead of doing channel-wise scaling:
@@
STE (ours) 1.64 80.7
XNOR-Net 1.64 79.4
Dorefa-Net 1.64 79.8
The experimental results of using different methods to deal with the sign function are shown in Tab. 4.3. The original STE method performs best among different methods. This conclusion is different from that in BNNs. We analyze that this is because both XNOR-Net (Rastegari et al., 2016) and Dorefa-Net (Zhou et al., 2016) modify the forward pass of sign function by adding per-channel scales or a constant scale, which contaminate the direction group and make it unable to focus on giving the correct binary outputs.
@@
Specifically, we keep the total trainable parameters unchanged by setting $r_{1}+r_{2}=2r$, and adjust the ratio of parameters in the magnitude group and direction group, which are controlled by $r_{1}$ and $r_{2}$, respectively. We conduct experiments on the commonsense reasoning dataset with LLaMA3-8B as our baseline model and set 15 different ratios. Other training hyper-parameters are kept the same as in the previous experiments. The results are shown in Fig. 2. We can see that the proposed Dual LoRA can consistently outperform LoRA, DoRA, and HiRA on a wide range of parameter ratio between the magnitude group and direction group (from 6% to 94%), which shows the robustness of our method.
@@ 6 Conclusion
7 Limitations
Our Dual LoRA achieves state-of-the-art performance on several benchmarks and baseline models. However, it incurs a slightly higher GPU memory and training time overhead compared to the original LoRA (less than 5%), due to the intermediate feature maps generated by the magnitude group and the direction group.
@@ Appendix A Experiments on Neural Language Generation (NLG)
Datasets. We conduct experiments on E2E NLG Challenge (Novikova et al., 2017) dataset. The E2E dataset consists of approximately 42,000 training data, 4,600 validation data, and 4,600 test data. Each sample is composed of a sequence of slot-value pairs $(x,y)$ and a corresponding natural language reference text.
Results. As in previous experiments, we reduce our rank to half of the other LoRA-based methods to ensure the same number of trainable parameters. All other hyper-parameters are the same as LoRA, except that we tune the learning rate. As the results shown in Tab. 4, Dual LoRA can consistently outperform all other competitors with baseline model GPT2_M and GPT2_L.
@@ B.1 Different forms of the update matrix
Dual LoRA ($r=16$) 0.83 71.3 83.4 81.0 88.1 83.7 85.0 71.0 82.4 80.7
Setting 1 ($r=32$) 1.64 69.3 83.0 79.6 87.6 82.8 85.9 69.3 82.2 80.0
Setting 2 ($r=32$) 1.64 66.7 82.0 79.8 86.2 82.0 84.4 68.3 82.6 79.0
Setting 3 ($r=32$) 0.84 62.0 48.4 33.2 7.9 0 24.1 25.4 28.0 28.6
Setting 3 ($r=64$) 1.64 62.2 49.9 33.5 25.6 50.5 25.6 26.6 25.8 37.5
@@
ReLU (ours) 71.3 83.4 81.0 88.1 83.7 85.0 71.0 82.4 80.7
Abs 66.1 79.7 77.5 83.2 79.2 80.7 66.5 78.2 76.4
Sigmoid 60.7 80.6 76.6 81.8 76.6 80.8 64.2 75.0 74.5